# Rémi A. Chou and Matthieu R. Bloch

### v1

@@ IV-B Single-Bidder Multi-Verifier Model and Results
    Reveal phase: Suppose that some verifiers may have dropped out of the network after the commit phase and $\mathcal{A}\subset\mathcal{B}$ is the set of available verifiers with $|\mathcal{A}|\geqslant 1$. The bidder reveals $(a,S)$ and chooses a verifier $b^{\star}\in\mathcal{A}$. Verifier $b^{\star}$ performs a test $\beta(V_{b^{\star}},a,S)$ that returns $1$ if the sequence $a$ is accepted and 00 otherwise.
@@
A rate $R$ is achievable if there exists a sequence of $(2^{n{R}},n)$ commitment schemes such that for any $\tilde{S}\in\mathcal{R}$, $a,a^{\prime}\in\mathcal{A}$ such that $a\neq a^{\prime}$,
[…]
@@
The correctness, concealment, and bindingness requirements in Definition 3 can be interpreted similarly to those in the settings discussed in Section IV-A.
@@ V-A Achievability scheme
The achievability proofs for the two theorems are similar, differing only in the set of allowed input distributions to the channel. To simultaneously capture both proofs, we define
$\bar{\mathcal{P}}(\mathcal{X}_{\mathcal{L}})\triangleq\begin{cases}\mathcal{P}^{\textup{I}}(\mathcal{X}_{\mathcal{L}})&\text{if the biddersare non-colluding}\\ \mathcal{P}(\mathcal{X}_{\mathcal{L}})&\text{if the bidders are colluding}\end{cases}.$
@@ V-A1 Definitions
The following notions of typicality will prove useful. Let $\epsilon>0$. For $x^{n}_{\mathcal{L}}\in\mathcal{X}_{\mathcal{L}}^{n}$, define
@@
[…]
Define also
@@
[…]
@@ V-A3 Concealment
$I(A_{\mathcal{L}};V)$ =$\stackrel{{\scriptstyle(a)}}{{=}}I(A_{\mathcal{L}};E_{\mathcal{L}})+I(A_{\mathcal{L}};V^{\prime}|E_{\mathcal{L}})$
@@
=$\stackrel{{\scriptstyle(b)}}{{=}}I(A_{\mathcal{L}};E_{\mathcal{L}})+I(A_{\mathcal{L}}K_{\mathcal{L}};V^{\prime})$
@@
=$\stackrel{{\scriptstyle(c)}}{{=}}I(A_{\mathcal{L}};E_{\mathcal{L}})+I(K_{\mathcal{L}};V^{\prime})$
@@
=$\stackrel{{\scriptstyle(e)}}{{=}}r_{\mathcal{L}}-H(K_{\mathcal{L}})+I(K_{\mathcal{L}};V^{\prime})$
@@
=$\stackrel{{\scriptstyle(f)}}{{=}}r_{\mathcal{L}}-H(K_{\mathcal{L}})+I(K_{\mathcal{L}};F_{\mathcal{L}}\bar{Y}^{n}),$ (40)
@@
Then, we upper bound the right-hand side of (40) using the version of the leftover hash lemma in Lemma 1, and we lower bound the min-entropies [18] appearing in Lemma 1 using Lemma 2.
Lemma 1 (Distributed leftover hash lemma, e.g., [19, Lemma 3]).
Consider a sub-normalized non-negative function $p_{X_{\mathcal{L}}Z}$ defined over ×$\bigtimes_{\ell\in\mathcal{L}}\mathcal{X}_{l}\times\mathcal{Z}$, where $X_{\mathcal{L}}\triangleq(X_{\ell})_{\ell\in\mathcal{L}}$ and, $\mathcal{Z}$, $\mathcal{X}_{l}$, $\ell\in\mathcal{L}$, are finite alphabets. For $\ell\in\mathcal{L}$, let $F_{\ell}:\{0,1\}^{n_{\ell}}\longrightarrow\{0,1\}^{r_{\ell}}$, be uniformly chosen in a family $\mathcal{F}_{\ell}$ of two-universal hash functions. For any $\mathcal{T}\subseteq\mathcal{L}$, define $r_{\mathcal{T}}\triangleq\sum_{i\in\mathcal{T}}r_{i}$. Define also ${F}_{\mathcal{L}}\triangleq(F_{\ell})_{\ell\in\mathcal{L}}$ and $F_{\mathcal{L}}(X_{\mathcal{L}})\triangleq\left(F_{\ell}(X_{\ell})\right)_{{\ell}\in\mathcal{L}}$. Then, for any $q_{Z}$ defined over $\mathcal{Z}$ such that $\textup{supp}(q_{Z})\subseteq\textup{supp}(p_{Z})$, we have
@@
$\stackrel{{\scriptstyle(c)}}{{\leqslant}}\epsilon+\mathbb{V}(w_{K_{\mathcal{L}}F_{\mathcal{L}}\bar{Y}^{n}},p_{U_{\mathcal{K}}}p_{U_{\mathcal{F}}}w_{\bar{Y}^{n}})+\mathbb{V}(p_{U_{\mathcal{K}}}p_{U_{\mathcal{F}}}w_{\bar{Y}^{n}},p_{U_{\mathcal{K}}}p_{U_{\mathcal{F}}}q_{\bar{Y}^{n}})$
@@
We conclude from (43) and [20, Lemma 2.7] that concealment holds with $(r_{l})_{{\ell}\in\mathcal{L}}$ such that for any $\mathcal{T}\subseteq\mathcal{L}$, $\lim_{n\to\infty}\frac{r_{\mathcal{T}}}{n}\leqslant H(X_{\mathcal{T}}|Y)-\delta$, $\delta>0$.
@@ V-A5 Achievable region and sum-rate
$=H(X_{\mathcal{U}}|Y)+H(X_{\mathcal{V}\backslash\mathcal{U}}|YX_{\mathcal{U}})+H(X_{\mathcal{V}}|Y)-H(X_{\mathcal{V}\backslash\mathcal{U}}|YX_{\mathcal{U}\cap\mathcal{V}})$
@@ V-B Converse
=$\stackrel{{\scriptstyle(a)}}{{=}}I(A_{\ell}S_{\ell};Y^{n}S^{\prime}_{\ell}|\bar{M}_{l,1:n-1}\bar{M}_{l,n,1:r_{n}-1}M_{l,n,r_{n}}X^{n}_{\ell})$
@@
=$\stackrel{{\scriptstyle(b)}}{{=}}I(A_{\ell}S_{\ell};Y^{n}S^{\prime}_{\ell}|\bar{M}_{l,1:n-1}\bar{M}_{l,n,1:r_{n}-1}X^{n}_{\ell})$ (45)
@@
=$\stackrel{{\scriptstyle(d)}}{{=}}I(A_{\ell}S_{\ell};Y^{n-1}S^{\prime}_{\ell}|\bar{M}_{l,1:n-1}X^{n}_{\ell})$
@@
=$\stackrel{{\scriptstyle(e)}}{{=}}I(A_{\ell}S_{\ell};Y^{n-1}S^{\prime}_{\ell}|\bar{M}_{l,1:n-1}X^{n-1}_{\ell})$ (46)
@@
    holds by repeating $n-1$ times the steps between (44) and (46).
@@
$F(x_{\ell}^{n},m_{\ell}|a_{\ell},s_{\ell})$ $\triangleq\mathbb{E}_{Y^{n}S^{\prime}_{\ell}|M_{\ell}=m_{\ell},X^{n}_{\ell}=x^{n}_{\ell},S_{\ell}=s_{\ell}}[\mathds{1}\{\beta_{\ell}(Y^{n},m_{\ell},S^{\prime}_{\ell},a_{\ell},s_{\ell})\}],$ (49)
@@
=$\stackrel{{\scriptstyle(a)}}{{=}}\mathbb{E}_{X_{\ell}^{n}M_{\ell}S_{\ell}|A_{\ell}=a_{\ell}}F(X_{\ell}^{n},M_{\ell}|a_{\ell})$
$\geqslant\sum_{s_{\ell}\in\mathcal{G}(a_{\ell})}p_{S_{\ell}|A_{\ell}=a_{\ell}}(s_{\ell})\mathbb{E}_{X_{\ell}^{n}M_{\ell}|A_{\ell}=a_{\ell},S_{\ell}=s_{\ell}}F(X_{\ell}^{n},M_{\ell}|a_{\ell})$
$\stackrel{{\scriptstyle(b)}}{{\geqslant}}\sum_{s_{\ell}\in\mathcal{G}(a_{\ell})}p_{S_{\ell}|A_{\ell}=a_{\ell}}(s_{\ell})\mathbb{E}_{X_{\ell}^{n}M_{\ell}|A_{\ell}=a_{\ell},S_{\ell}=s_{\ell}}F(X_{\ell}^{n},M_{\ell}|a_{\ell},s_{\ell})$
=$\stackrel{{\scriptstyle(c)}}{{=}}\sum_{s_{\ell}\in\mathcal{G}(a_{\ell})}p_{S_{\ell}|A_{\ell}=a_{\ell}}(s_{\ell})\mathbb{E}_{Y^{n}S^{\prime}_{\ell}M_{\ell}|S_{\ell}=s_{\ell},A_{\ell}=a_{\ell}}[\mathds{1}\{\beta(Y^{n},M_{\ell},S^{\prime}_{\ell},a_{\ell},s_{\ell})\}]$
=$\stackrel{{\scriptstyle(d)}}{{=}}\sum_{s_{\ell}\in\mathcal{G}(a_{\ell})}p_{S_{\ell}|A_{\ell}=a_{\ell}}(s_{\ell})f(a_{\ell},s_{\ell})$
>$\stackrel{{\scriptstyle(e)}}{{>}}\sum_{s_{\ell}\in\mathcal{G}(a_{\ell})}p_{S_{\ell}|A_{\ell}=a_{\ell}}(s_{\ell})(1-\gamma)$
@@
    =$\stackrel{{\scriptstyle(i)}}{{=}}\mathbb{E}_{X_{\ell}^{n}M_{\ell}|A_{\ell}=a_{\ell},S_{\ell}=s_{\ell}}\mathbb{E}_{Y^{n}S^{\prime}_{\ell}|M_{\ell}=M_{\ell},X^{n}_{\ell}=X^{n}_{\ell},S_{\ell}=s_{\ell}}[\mathds{1}\{\beta_{\ell}(Y^{n},M_{\ell},S^{\prime}_{\ell},a_{\ell}^{\dagger},s_{\ell}^{\dagger})\}]$
    =$\stackrel{{\scriptstyle(ii)}}{{=}}\mathbb{E}_{X_{\ell}^{n}M_{\ell}|A_{\ell}=a_{\ell},S_{\ell}=s_{\ell}}\mathbb{E}_{Y^{n}S^{\prime}_{\ell}|M_{\ell}=M_{\ell},X^{n}_{\ell}=X^{n}_{\ell},S_{\ell}=s_{\ell},A_{\ell}=a_{\ell}}[\mathds{1}\{\beta(Y^{n},M_{\ell},S^{\prime}_{\ell},a_{\ell}^{\dagger},s_{\ell}^{\dagger})\}]$
    $=\mathbb{E}_{X_{\ell}^{n}Y^{n}S^{\prime}_{\ell}M_{\ell}|S_{\ell}=s_{\ell},A_{\ell}=a_{\ell}}[\mathds{1}\{\beta(Y^{n},M_{\ell},S^{\prime}_{\ell},a_{\ell}^{\dagger},s_{\ell}^{\dagger})\}]$
    =$\stackrel{{\scriptstyle(iii)}}{{=}}\mathbb{E}_{Y^{n}S^{\prime}_{\ell}M_{\ell}|S_{\ell}=s_{\ell},A_{\ell}=a_{\ell}}[\mathds{1}\{\beta(Y^{n},M_{\ell},S^{\prime}_{\ell},a_{\ell}^{\dagger},s_{\ell}^{\dagger})\}],$ (54)
@@
      holds because $A_{\ell}-(M_{\ell},X^{n}_{\ell},S_{\ell})-(Y^{n},S^{\prime}_{\ell})$ forms a Markov chain since $I(A_{\ell};Y^{n}S^{\prime}_{\ell}|M_{\ell}X^{n}_{\ell}S_{\ell})\leqslant I(A_{\ell}S_{\ell};Y^{n}S^{\prime}_{\ell}|M_{\ell}X^{n}_{\ell})=0$ by Lemma  4;
@@
=$\stackrel{{\scriptstyle(a)}}{{=}}\mathbb{E}_{X_{\ell}^{n}M_{\ell}|A_{\ell}=a_{\ell}}[{F}(X_{\ell}^{n},M_{\ell}|a_{\ell}^{*},s_{\ell}^{*})]$
=$\stackrel{{\scriptstyle(b)}}{{=}}\mathbb{E}_{S_{\ell}|A_{\ell}=a_{\ell}}\mathbb{E}_{X_{\ell}^{n}M_{\ell}|A_{\ell}=a_{\ell},S_{\ell}=S_{\ell}}[{F}(X_{\ell}^{n},M_{\ell}|a_{\ell}^{*},s_{\ell}^{*})]$
=$\stackrel{{\scriptstyle(c)}}{{=}}\mathbb{E}_{S_{\ell}|A_{\ell}=a_{\ell}}\mathbb{E}_{Y^{n}S^{\prime}_{\ell}M_{\ell}|S_{\ell}=S_{\ell},A_{\ell}=a_{\ell}}[\mathds{1}\{\beta(Y^{n},M_{\ell},S^{\prime}_{\ell},a_{\ell}^{*},s_{\ell}^{*})\}]$
@@
=$\stackrel{{\scriptstyle(c)}}{{=}}H(X^{n}_{\mathcal{T}}|Y^{n})$

### v2

@@ IV-B Single-Bidder Multi-Verifier Model and Results
    Reveal phase: Suppose that some verifiers may have dropped out of the network after the commit phase and $\mathcal{A}\subset\mathcal{B}$ is the set of available verifiers with $|\mathcal{A}|\geqslant 1$. The bidder reveals $(a,S)$ and chooses a verifier $b^{\star}\in\mathcal{A}$ at random. Verifier $b^{\star}$ performs a test $\beta(V_{b^{\star}},a,S)$ that returns $1$ if the sequence $a$ is accepted and 0 otherwise.
@@
A rate $R$ is achievable if there exists a sequence of $(2^{n{R}},n)$ commitment schemes such that for any $\tilde{S}\in\mathcal{R}$, $a,\tilde{a}\in\mathcal{A}$ such that $a\neq\tilde{a}$,
$\lim_{n\to\infty}\max_{b\in\mathcal{B}}\mathbb{P}[\beta(V_{b},a,S)=0]$ $=0,\text{ (correctness) }$ (37)
$\lim_{n\to\infty}\max_{b\in\mathcal{B}}I(A;V_{b})$ $=0,\text{ (concealement) }$ (38)
@@
The correctness, concealment, and bindingness requirements in Definition 3 can be interpreted similarly to those in the settings discussed in Section IV-A. Note that Definition 3 ensures that correctness, concealement, and bindingness are satisfied irrespective of $\mathcal{A}\subset\mathcal{B}$ such that $|\mathcal{A}|\geqslant 1$, and for any choice $b^{\star}\in\mathcal{A}$.
@@ V-A Achievability scheme
The achievability proof is inspired by the proof for point-to-point channel from [4] and relies on hash challenges, typicality, and privacy amplification [19], adapted to a distributed setting. The achievability proof holds for Theorem 1 and Theorem 2, differing only in the set of allowed input distributions to the channel. To simultaneously capture both theorems, we define
[…]
@@ V-A1 Definitions
The following notions of typicality will prove useful. Let $\epsilon>0$. For $x^{n}_{\mathcal{L}}\in\mathcal{X}_{\mathcal{L}}^{n}$, define $\mathcal{T}^{n}_{W,\epsilon}({x}_{\mathcal{L}}^{n})$ and $\mathcal{T}_{\epsilon}^{n}(q_{X_{\mathcal{L}}Y})$ in (40) and (41).
@@
 $\phantom{---}\left.\text{ and }W_{x_{\mathcal{L}}}(y)=0\implies\sum_{i=1}^{n}\frac{\mathds{1}\{(x_{\mathcal{L}},y)\!=\!(x_{\mathcal{L},i},y_{i})\}}{n}=0\right\}.$ (40)
@@
 $\phantom{---}\left.\text{ and }q_{X_{\mathcal{L}}Y}(x_{\mathcal{L}},y)=0\implies\sum_{i=1}^{n}\frac{\mathds{1}\{(x_{\mathcal{L}},y)\!=\!(x_{\mathcal{L},i},y_{i})\}}{n}=0\right\}.$ (41)
@@ V-A3 Concealment
$I(A_{\mathcal{L}};V)$ $\stackrel{{\scriptstyle(a)}}{{=}}I(A_{\mathcal{L}};E_{\mathcal{L}})+I(A_{\mathcal{L}};V^{\prime}|E_{\mathcal{L}})$
@@
$\stackrel{{\scriptstyle(b)}}{{=}}I(A_{\mathcal{L}};E_{\mathcal{L}})+I(A_{\mathcal{L}}K_{\mathcal{L}};V^{\prime})$
@@
$\stackrel{{\scriptstyle(c)}}{{=}}I(A_{\mathcal{L}};E_{\mathcal{L}})+I(K_{\mathcal{L}};V^{\prime})$
@@
$\stackrel{{\scriptstyle(e)}}{{=}}r_{\mathcal{L}}-H(K_{\mathcal{L}})+I(K_{\mathcal{L}};V^{\prime})$
@@
$\stackrel{{\scriptstyle(f)}}{{=}}r_{\mathcal{L}}-H(K_{\mathcal{L}})+I(K_{\mathcal{L}};F_{\mathcal{L}}\bar{Y}^{n}),$ (42)
@@
Then, we upper bound the right-hand side of (42) using the version of the distributed leftover hash lemma [21] in Lemma 1, and we lower bound the min-entropies [22, 23] appearing in Lemma 1 using Lemma 2.
Lemma 1 (Distributed leftover hash lemma, e.g., [24, Lemma 3]).
Consider a sub-normalized non-negative function $p_{X_{\mathcal{L}}Z}$ defined over $\bigtimes_{\ell\in\mathcal{L}}\mathcal{X}_{l}\times\mathcal{Z}$, where $X_{\mathcal{L}}\triangleq(X_{\ell})_{\ell\in\mathcal{L}}$ and, $\mathcal{Z}$, $\mathcal{X}_{l}$, $\ell\in\mathcal{L}$, are finite alphabets. For $\ell\in\mathcal{L}$, let $F_{\ell}:\{0,1\}^{n_{\ell}}\longrightarrow\{0,1\}^{r_{\ell}}$, be uniformly chosen in a family $\mathcal{F}_{\ell}$ of two-universal hash functions. For any $\mathcal{T}\subseteq\mathcal{L}$, define $r_{\mathcal{T}}\triangleq\sum_{i\in\mathcal{T}}r_{i}$. Define also ${F}_{\mathcal{L}}\triangleq(F_{\ell})_{\ell\in\mathcal{L}}$ and $F_{\mathcal{L}}(X_{\mathcal{L}})\triangleq\left(F_{\ell}(X_{\ell})\right)_{{\ell}\in\mathcal{L}}$. Then, for any $q_{Z}$ defined over $\mathcal{Z}$ such that $\textup{supp}(q_{Z})\subseteq\textup{supp}(p_{Z})$, we have
@@
$\stackrel{{\scriptstyle(c)}}{{\leqslant}}\epsilon+\mathbb{V}(w_{K_{\mathcal{L}}F_{\mathcal{L}}\bar{Y}^{n}},p_{U_{\mathcal{K}}}p_{U_{\mathcal{F}}}w_{\bar{Y}^{n}})$
$\phantom{---}+\mathbb{V}(p_{U_{\mathcal{K}}}p_{U_{\mathcal{F}}}w_{\bar{Y}^{n}},p_{U_{\mathcal{K}}}p_{U_{\mathcal{F}}}q_{\bar{Y}^{n}})$
@@
Finally, we conclude by invoking the following lemma.
Lemma 3 ([25]).
Let $X$ and $Y$ be discrete random variables defined over $\mathcal{X}$ and $\mathcal{Y}$, respectively, with joint distribution $p_{XY}$. If $|\mathcal{X}|\geqslant 4$, then
$I(X;Y)\leqslant\mathbb{V}(p_{XY},p_{X}p_{Y})\log\frac{|\mathcal{X}|}{\mathbb{V}(p_{XY},p_{X}p_{Y})}.$
By (45) and Lemma 3, concealment holds with $(r_{l})_{{\ell}\in\mathcal{L}}$ such that for any $\mathcal{T}\subseteq\mathcal{L}$, $\lim_{n\to\infty}\frac{r_{\mathcal{T}}}{n}\leqslant H(X_{\mathcal{T}}|Y)-\delta$, $\delta>0$.
@@ V-A4 Bindingness
A close inspection of the proof reveals that the lemma holds regardless of whether the joint distribution of the inputs is a product of the marginals or not.
@@ V-A5 Achievable region and sum-rate
$=H(X_{\mathcal{U}}|Y)+H(X_{\mathcal{V}\backslash\mathcal{U}}|YX_{\mathcal{U}})$
$\phantom{--}+H(X_{\mathcal{V}}|Y)-H(X_{\mathcal{V}\backslash\mathcal{U}}|YX_{\mathcal{U}\cap\mathcal{V}})$
@@ V-B Converse
$\stackrel{{\scriptstyle(a)}}{{=}}I(A_{\ell}S_{\ell};Y^{n}S^{\prime}_{\ell}|\bar{M}_{l,1:n-1}\bar{M}_{l,n,1:r_{n}-1}M_{l,n,r_{n}}X^{n}_{\ell})$
@@
$\stackrel{{\scriptstyle(b)}}{{=}}I(A_{\ell}S_{\ell};Y^{n}S^{\prime}_{\ell}|\bar{M}_{l,1:n-1}\bar{M}_{l,n,1:r_{n}-1}X^{n}_{\ell})$ (47)
@@
$\stackrel{{\scriptstyle(d)}}{{=}}I(A_{\ell}S_{\ell};Y^{n-1}S^{\prime}_{\ell}|\bar{M}_{l,1:n-1}X^{n}_{\ell})$
@@
$\stackrel{{\scriptstyle(e)}}{{=}}I(A_{\ell}S_{\ell};Y^{n-1}S^{\prime}_{\ell}|\bar{M}_{l,1:n-1}X^{n-1}_{\ell})$ (48)
@@
    holds by repeating n1n-1 times the steps between (46) and (48).
@@
$F(x_{\ell}^{n},m_{\ell}|a_{\ell},s_{\ell})$
$\triangleq\mathbb{E}_{Y^{n}S^{\prime}_{\ell}|M_{\ell}=m_{\ell},X^{n}_{\ell}=x^{n}_{\ell},S_{\ell}=s_{\ell}}[\mathds{1}\{\beta_{\ell}(Y^{n},m_{\ell},S^{\prime}_{\ell},a_{\ell},s_{\ell})\}],$ (51)
@@
$\stackrel{{\scriptstyle(a)}}{{=}}\mathbb{E}_{X_{\ell}^{n}M_{\ell}S_{\ell}|A_{\ell}=a_{\ell}}F(X_{\ell}^{n},M_{\ell}|a_{\ell})$
$\geqslant\textstyle\sum_{s_{\ell}\in\mathcal{G}(a_{\ell})}p_{S_{\ell}|A_{\ell}=a_{\ell}}(s_{\ell})$
$\phantom{----}\times\mathbb{E}_{X_{\ell}^{n}M_{\ell}|A_{\ell}=a_{\ell},S_{\ell}=s_{\ell}}F(X_{\ell}^{n},M_{\ell}|a_{\ell})$
$\stackrel{{\scriptstyle(b)}}{{\geqslant}}\textstyle\sum_{s_{\ell}\in\mathcal{G}(a_{\ell})}p_{S_{\ell}|A_{\ell}=a_{\ell}}(s_{\ell})$
$\phantom{----}\times\mathbb{E}_{X_{\ell}^{n}M_{\ell}|A_{\ell}=a_{\ell},S_{\ell}=s_{\ell}}F(X_{\ell}^{n},M_{\ell}|a_{\ell},s_{\ell})$
$\stackrel{{\scriptstyle(c)}}{{=}}\textstyle\sum_{s_{\ell}\in\mathcal{G}(a_{\ell})}p_{S_{\ell}|A_{\ell}=a_{\ell}}(s_{\ell})$
$\phantom{----}\times\mathbb{E}_{Y^{n}S^{\prime}_{\ell}M_{\ell}|S_{\ell}=s_{\ell},A_{\ell}=a_{\ell}}[\mathds{1}\{\beta(Y^{n},M_{\ell},S^{\prime}_{\ell},a_{\ell},s_{\ell})\}]$
$\stackrel{{\scriptstyle(d)}}{{=}}\textstyle\sum_{s_{\ell}\in\mathcal{G}(a_{\ell})}p_{S_{\ell}|A_{\ell}=a_{\ell}}(s_{\ell})f(a_{\ell},s_{\ell})$
$\stackrel{{\scriptstyle(e)}}{{>}}\textstyle\sum_{s_{\ell}\in\mathcal{G}(a_{\ell})}p_{S_{\ell}|A_{\ell}=a_{\ell}}(s_{\ell})(1-\gamma)$
@@
    $\!\!\!\!\!\stackrel{{\scriptstyle(i)}}{{=}}\mathbb{E}_{X_{\ell}^{n}M_{\ell}|A_{\ell}=a_{\ell},S_{\ell}=s_{\ell}}\mathbb{E}_{Y^{n}S^{\prime}_{\ell}|M_{\ell}=M_{\ell},X^{n}_{\ell}=X^{n}_{\ell},S_{\ell}=s_{\ell}}$
    $\phantom{-------l-----}[\mathds{1}\{\beta_{\ell}(Y^{n},M_{\ell},S^{\prime}_{\ell},a_{\ell}^{\dagger},s_{\ell}^{\dagger})\}]$
    $\!\!\!\!\!\!\stackrel{{\scriptstyle(ii)}}{{=}}\mathbb{E}_{X_{\ell}^{n}M_{\ell}|A_{\ell}=a_{\ell},S_{\ell}=s_{\ell}}\mathbb{E}_{Y^{n}S^{\prime}_{\ell}|M_{\ell}=M_{\ell},X^{n}_{\ell}=X^{n}_{\ell},S_{\ell}=s_{\ell},A_{\ell}=a_{\ell}}$
    $\phantom{-------l-----}[\mathds{1}\{\beta_{l}(Y^{n},M_{\ell},S^{\prime}_{\ell},a_{\ell}^{\dagger},s_{\ell}^{\dagger})\}]$
    $\!\!\!\!\!=\mathbb{E}_{X_{\ell}^{n}Y^{n}S^{\prime}_{\ell}M_{\ell}|S_{\ell}=s_{\ell},A_{\ell}=a_{\ell}}[\mathds{1}\{\beta_{l}(Y^{n},M_{\ell},S^{\prime}_{\ell},a_{\ell}^{\dagger},s_{\ell}^{\dagger})\}]$
    $\!\!\!\!\!\!\!\stackrel{{\scriptstyle(iii)}}{{=}}\mathbb{E}_{Y^{n}S^{\prime}_{\ell}M_{\ell}|S_{\ell}=s_{\ell},A_{\ell}=a_{\ell}}[\mathds{1}\{\beta_{l}(Y^{n},M_{\ell},S^{\prime}_{\ell},a_{\ell}^{\dagger},s_{\ell}^{\dagger})\}],$ (56)
@@
      holds because $A_{\ell}-(M_{\ell},X^{n}_{\ell},S_{\ell})-(Y^{n},S^{\prime}_{\ell})$ forms a Markov chain since $I(A_{\ell};Y^{n}S^{\prime}_{\ell}|M_{\ell}X^{n}_{\ell}S_{\ell})\leqslant I(A_{\ell}S_{\ell};Y^{n}S^{\prime}_{\ell}|M_{\ell}X^{n}_{\ell})=0$ by Lemma  5;
@@
$\stackrel{{\scriptstyle(a)}}{{=}}\mathbb{E}_{X_{\ell}^{n}M_{\ell}|A_{\ell}=a_{\ell}}[{F}(X_{\ell}^{n},M_{\ell}|a_{\ell}^{*},s_{\ell}^{*})]$
$\stackrel{{\scriptstyle(b)}}{{=}}\mathbb{E}_{S_{\ell}|A_{\ell}=a_{\ell}}\mathbb{E}_{X_{\ell}^{n}M_{\ell}|A_{\ell}=a_{\ell},S_{\ell}=S_{\ell}}[{F}(X_{\ell}^{n},M_{\ell}|a_{\ell}^{*},s_{\ell}^{*})]$
$\stackrel{{\scriptstyle(c)}}{{=}}\mathbb{E}_{S_{\ell}|A_{\ell}=a_{\ell}}\mathbb{E}_{Y^{n}S^{\prime}_{\ell}M_{\ell}|S_{\ell}=S_{\ell},A_{\ell}=a_{\ell}}[\mathds{1}\{\beta_{l}(Y^{n}\!\!,M_{\ell},S^{\prime}_{\ell},a_{\ell}^{*},s_{\ell}^{*})\!\}]$
@@
$\stackrel{{\scriptstyle(c)}}{{=}}H(X^{n}_{\mathcal{T}}|Y^{n})$